This quation in differential equations. i need in in clear and short hand writing please. thansk
1. Solve the following IVP. z?y = 1-2xy y(1) = 4 / (1) = 2
Answer #1

Deve zy = 1-2 ay y(l)=2 22y tray= 1 0 This is Cauchys DE put ace? . Z=lud. 224-D(0-1) ny = Dy. put these values in eqAD-y +Ay = 1 Amilay et sm-1) + 2m 50. mEm +2m=0 mtm=0 m (m +1)=0 m=0,-| -lud. Y Ct4e-2 Yecteel = c+qe Yo = citezat TYCE Cpt

If you have any questions please let me know

Thank you

Please give me thumb up..